OpenStudy (anonymous):

If cscx= (4√3)/3 and cosx<0, find sin x

OpenStudy (anonymous):

Using trig definitions and identities: 1. sin x = 1/csc x = 1/3 2. cos²x = 1 - sin²x = 1 - 1/9 = 8/9. Since cos x < 0 we take the negative square root: cos x = -(√8)/3 = -2(√2)/3. 3. tan x = sin x/cos x = (1/3) / -2(√2)/3 = -1/(2√2) = -(√2)/4. 4. cot x = 1/tan x = -2√2. 5. sec x = 1/cos x = -3/(2√2) = -3(√2)/4

OpenStudy (anonymous):

metal